OpenStudy (anonymous):

how do you know when to try a trig substitution for an integral? sometimes it doesn't seem obvious :/

sam (.sam.):

square root with both square terms, if you know what i mean :)

OpenStudy (marco26):

you can use them when the integrand contains radical form (only second radical) of \[\sqrt{a^2-u^2}\]\[\sqrt{a^2+u^2}\]\[\sqrt{u^2-a^2}\] or even without the radical, just follow the pattern. a is constant

OpenStudy (anonymous):

when the integrand involves 1.\[a^2 -x^2\] 2.\[a^2 + x^2\] 3.\[x^2 -a^2\] try 1. \[x=asin \theta\] 2. \[x=atan\] 3.\[x=asec \theta\]

sam (.sam.):

2. \[x=atan \theta\]

OpenStudy (anonymous):

@.Sam. yup

OpenStudy (anonymous):

sometimes we have the second radical raised to a power, and that works to eg \[\int\limits{\frac{1}{(x^2+a^2)^{3/2}}}dx\]

OpenStudy (anonymous):

so should i think of that as \[(\sqrt{x^2+a^2})^3\] i presume

OpenStudy (anonymous):

will it work for any f(x^2 +a^2) ?

OpenStudy (anonymous):

\[\int\limits{f(x^2 + a^2)}dx\] let x = atanu \[\int\limits{f(\sec^2u)}\sec^2u dx\] so i guess we need f such that it cancels somehow. in the previous comment it comes to integral(cosu du) sinu + C but u = arctan(x/a) so its sin(arctan(x/a)) ? can this be written without trig ?

OpenStudy (blockcolder):

Use right triangles.